# Quantum State Generation with Structure-Preserving Diffusion Model

### v1

@@ Abstract
This article considers the generative modeling of the states of quantum systems, and an approach based on denoising diffusion model is proposed. The key contribution is an algorithmic innovation that respects the physical nature of quantum states. More precisely, the commonly used density matrix representation of mixed-state has to be complex-valued Hermitian, positive semi-definite, and trace one. Generic diffusion models, or other generative methods, may not be able to generate data that strictly satisfy these structural constraints, even if all training data do. To develop a machine learning algorithm that has physics hard-wired in, we leverage the recent development of Mirror Diffusion Model and design a previously-unconsidered mirror map, to enable strict structure-preserving generation. Both unconditional generation and conditional generation via classifier-free guidance are experimentally demonstrated efficacious, the latter even enabling the design of new quantum states when generated on unseen labels.
quantum density matrix, mixed state synthesis, diffusion generative modeling, constrained generation
@@
1 Introduction
Generative modeling, a specialized subfield of machine learning, is dedicated to creating models capable of generating new data points that closely match the statistical distribution of a given dataset. These models adeptly capture and replicate the patterns, structures, and characteristics of the input data, allowing them to yield new instances that seem to originate from the same distribution. Among them, diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) has risen to prominence for its remarkable ability to produce high-quality data, especially notable in image generation (Rombach et al., 2022; Ho et al., 2022). The fundamental mechanism of diffusion model starts with adding noise to the data and progressively altering it into a tractable Gaussian prior with analytically available transition kernels. The model then effectively reverses this noise addition, reconstructing the data from its noised state through the application of learned score functions (Anderson, 1982; Haussmann & Pardoux, 1986). This process not only results in easily obtainable conditional score functions but also simplifies regression objectives, significantly boosting their capability in managing complex, high-dimensional data distributions.
Diffusion model has showcased impressive capabilities in generating various types of dataset across different application area, such as video synthesis (Ge et al., 2023; Blattmann et al., 2023), language modeling (Austin et al., 2021; Lou et al., 2023) and point cloud generation (Zeng et al., 2022; Luo & Hu, 2021). With the recent rise of AI4Science, the potential of diffusion model in tackling scientific problems also started being explored, including applications in biology (Jumper et al., 2021; Watson et al., 2022), chemistry (Duan et al., 2023; Hoogeboom et al., 2022), and climate science (Mardani et al., 2024). Other important notions relevant to this article include constrained generations (e.g., Liu et al., 2023; Fishman et al., 2023b, a; Lou & Ermon, 2023) and conditional generations (e.g., Dhariwal & Nichol, 2021; Ho & Salimans, 2022).
@@ 1 Introduction
What is under-studied but important is the (classical) generative modeling of quantum data. This is not the same as quantum computing or quantum machine learning, because the goal is not to design quantum algorithms, but to develop algorithms that run on classical computers for handling quantum data. Similar to the already extremely successful application of the generation (and thus design) of molecular configurations (e.g., Watson et al., 2023; Duan et al., 2023), it would be useful to generate, for example, the states of quantum systems, before quantum computers become accessible to general researchers.
@@
To be able to represent quantum states, which in general can be mixed states, one needs to be mindful of their structural constraints that originated from physics. For example, it is standard in quantum information and quantum computing to represent mixed states as complex-valued density matrices (Nielsen & Chuang, 2010), but such matrices have to be Hermitian, positive semidefinite, and trace one. Data-based generation of quantum density matrices therefore amounts to generative modeling in non-Euclidean spaces. Without strictly satisfying the structural constraints, the generated matrices will no longer correspond to quantum states (e.g., Huang et al., 2020). Unfortunately, generic generative modeling approaches will almost surely violate the constraints due to indispensable errors originated from finite data and compute. This article proposes the first machine learning methodology that generates density matrices similar to training data and strictly satisfying their structural constraints, based on hard-wiring these constraints into diffusion models via technique from convex optimization.
@@
Imperative to mention is, machine learning has already been demonstrated promising for quantum problems, and it is rapidly becoming popular in the domain. There is even an interesting research frontier on quantum algorithms for generative modeling (e.g., Dallaire-Demers & Killoran, 2018; Khoshaman et al., 2018; Chakrabarti et al., 2019; Stein et al., 2021; Zhang et al., 2024; Chen & Zhao, 2024); however, this article still focuses on classical algorithm as quantum computer is not yet widely available, and existing results did not generate mixed states larger than 3-qubits. Meanwhile, classical algorithms have already been considered for many important quantum problems, such as the classification and regression of quantum states (e.g., Huang et al., 2020, 2022) and quantum state tomography (e.g., Torlai et al., 2018; Torlai & Melko, 2018; Carrasquilla et al., 2019; Vicentini et al., 2022; Jayakumar et al., 2023). The latter, roughly speaking, corresponds to the process of estimating one latent quantum density matrix through repeated, but necessarily noisy measurements of many copies of the same state. It is very important to clarify the similarity and difference between this work and quantum state tomography:
@@
In fact, some may feel that quantum state tomography is already generative modeling, in a broad sense, because once the latent deterministic density matrix is recovered from noisy measurements, one can produce additional, synthetic classical measurements, which are random and ideally identically distributed with the training data. However, it is actually very different from the generative modeling of quantum states considered here, which might be in a narrow sense, however being a common setup in (classical) machine learning. More precisely, what this article considers is a collection of many different quantum mixed states, instead of classical observations of a single quantum state, and the goal is to generate more mixed states that are similar to the given ones. The problem of using machine learning to infer a single quantum state from classical observational data, i.e. state tomography, has already been considered by a series of seminal work mentioned above; there the probability distribution of data is uniquely determined by a latent quantum state, and thus one essentially needs to learn a $D$-dimensional vector (if the latent state is an $N$-qubit mixed state represented by a $2^{N}\times 2^{N}$ density matrix, then $D=2^{2N}$). The new task of state generative modeling in this article, on the other hand, assumes mixed states are already given and they follow some unknown latent probability distribution, which can be represented by a probability density function of $D$ variables, and the goal is in some sense to learn this function, which is exponentially more complex than a $D$-dimensional vector. An additional difference between this work and existing milestones on generative modeling for state tomography, from a machine learning methodology point of view, is that this work is based on denoising diffusion for generative modeling, while work cited above used, e.g., EBM, RBM, RNN instead.
@@
One perspective is to actually create innovative states. Unconditional generation will simply produce states similar to those in the training data, but modern (diffusion) generative modeling technique also unleashes the ability of conditional generation. More precisely, given training data that carry different labels (e.g., traditionally labels could be ‘astronaut’, ‘horse’, etc., but in our case they can be ‘insulating’, ‘magnetic’, ‘topological’, etc.), we can not only generate under a label that already has training data (e.g., ‘astronaut’, or ‘magnetic’), but also generate under a label that has no training data (e.g., the famous ‘horse-riding astronaut’ (Heaven, 2022), or, this time, a motivating perspective of creating ‘topological magnetic’ states). This article humbly starts a first step toward this goal of scientific innovation, essentially via the interpolation/extrapolation of quantum data. The inter-/extra-polation, however, is in the sense of probability distributions but not sample points, because when one tries to design new states, it may not make sense to take, e.g., a specific ‘topological’ state and a specific ‘magnetic’ state and force a new ‘topological magnetic’ state based on them two alone. Instead, our conditional generative modeling approach hybrids multiple probability distributions and innovate just like prevailing generative AI tools.
@@
Another application is, generative modeling can also estimate the probability distribution of states in the data, hence providing further quantitative insights for probing the physical process behind it. Some consideration of this aspect was already included in the state tomography literature (e.g., Carrasquilla et al., 2019), but not for an ensemble of states. Moreover, diffusion model excels in this regard. As a more recent generative approach, it is posited to be particularly effective for estimating density near the data manifold, even when data is high-dimensional but sample size is small, as long as there is a low-dimensional structure in the data. This is oftentimes the case for image and video generations, and their empirical successes were the root of this belief. Quantum density matrices data could be similar; at least, they are high-dimensional since dimension grows exponentially with the number of qubits, but scarce due to the cost of experiments. This is a case where diffusion generative model could be advantageous over traditional approaches.
@@
In addition, generative model can serve as a sampler for crafting synthetic data that are experimentally expensive to harvest. Synthetic data that are statistically similar to training data are useful for discovering system properties through helping estimate observable values. The seminal work (Carrasquilla et al., 2019) showcases one such application for state-tomography type generative modeling. The idea is, generative modeling could help obtain more virtual ‘measurements’ from physical ones, hence improving the accuracy of the estimated latent state. Analogously, generative modeling of a distribution of mixed states can help the investigation of stochastic quantum systems, such as Random Transverse-field Ising Models (RTFIM) (Fisher, 1992, 1995; Kovács, 2022) and measurement-altered quantum critical systems (Murciano et al., 2023). The generated states can empower more accurate estimation of random parameters in the system, such as coupling and transverse field strengths.
@@ 2.1 Unconditional Generation via Diffusion Model
Denoising diffusion generative models in Euclidean space admit many descriptions, with focuses on different perspectives, and here we adopt the stochastic differential equations (SDE) description (Song et al., 2021). Given samples of $\mathbb{R}^{d}$-valued random variable $x_{0}$ that follows the data distribution $p_{0}$, denoising diffusion adopts a forward noising process followed by a backward denoising generation process to generate more samples of $p_{0}$.
@@
The backward process then utilizes the time-reversal of the SDE (1) (Anderson, 1982). More precisely, if one considers
$\mathrm{d}y_{\tau}$ $=(-f(y_{\tau},T-\tau)+g(T-\tau)^{2}s(y_{\tau},T-\tau))\mathrm{d}\tau$
$+g(T-\tau)\mathrm{d}w_{\tau},$ (2)
with $y_{0}\sim p_{T}$, where $s$, known as the score function, is given by
$s(x,t)=\nabla\log p_{t}(x),$
then we have $y_{\tau}\sim p_{T-\tau}$, i.e. $y_{\tau}=x_{T-\tau}$ in distribution. In particular, the $T$-time evolution of (2.1), $y_{T}$, will follow the data distribution $p_{0}$.
@@
To do so, the score $s(x,t)=\nabla\log p_{t}(x)$ is often approximated using a neural network $s_{\theta}(x,t)$. For linear forward SDE, it is typically trained by minimizing an objective based on denoising score matching (Vincent, 2011), namely
@@
where $\nabla\log p_{t}(x_{t}|x_{0})$ is the conditional score derived from the solution of (1) with a given initial condition.
@@ 2.2 Diffusion Guidance for Conditional Generation
Dhariwal & Nichol (2021) exploited this fact and proposed the technique of classifier-based diffusion guidance, originally for boosting the sample quality generated by diffusion models. Instead of $p(x|y)$, they proposed to sample from a condition-enhanced distribution,
$p_{\gamma}(x|y)\propto p(x)p(y|x)^{\gamma}.$
The score of this condition-enhanced distribution can be computed by
@@
$\gamma$ is called the strength of the guidance, which amplifies the influence of the conditioning when setting to a scale larger than $1$. When $\gamma>1$, the distribution is sharpened and focused onto its mode that corresponds to the condition $y$ (Dieleman, 2022). By tuning $\gamma$, classifier guidance allows us to capture the influence of the condition signal.
However, classifier-based diffusion guidance can become impractical due to the expensive cost of training a separate classifier model. Such a classifier often requires to be noise-robust to handle the noise-corrupted input $x$ during the sampling process. Ho & Salimans (2022) proposed classifier-free diffusion guidance to sample from the condition-enhanced distribution $p_{\gamma}(x|y)$ without the need for an extra discriminative model. Their insight came from another writing of Bayes’ rule: $\nabla\log p_{\gamma}(x|y)$ can also be decomposed as,
@@ 2.3 Density Matrix Representation of Quantum State
Precisely, a (quantum) density matrix is a matrix with each element being a complex number, which satisfies the following structural constraints (e.g., Nielsen & Chuang, 2010):
@@
A $n-$qubit quantum state $\rho$ is an element of $\mathbb{C}^{+}_{2^{n}}$ with trace 1.
@@ 3 Method
Our task is to generate new samples of quantum mixed states, from a distribution characterized by the training data. Out of the three structure constraints of density matrix representation of quantum states, Hermitianity is easy to handle as one can just work with the upper triangular part of the matrix. Trace one condition will only be minimally violated due to generation error, and we simply normalize the generated result by dividing its trace to strictly enforce this constraint. The most nontrivial constraint is the positive semi-definiteness. Since this is a convex constraint, we can leverage the recent approach of Mirror Diffusion Model (MDM) (Liu et al., 2023), which is a class of diffusion models that generate data on convex-constrained sets, to enable exact generation on $\mathbb{C}^{+}_{n}$. To achieve this goal, we work out a complex and multivariate extension of the results in MDM.
@@ 3.1 Mirror Map with Real Numbers
Its inverse enjoys a pleasant characterization from convex analysis (Rockafellar, 2015). Namely, by defining the dual function of $\phi$ as $\phi^{*}(y)=\sup_{x\in\mathcal{M}}\langle x,y\rangle-\phi(x),$ the inverse (as a mapping) of $\nabla\phi$ is the same as the gradient of $\phi^{*}$, i.e.
$\nabla\phi^{*}(\nabla\phi(x))=x,\quad\forall x\in\mathcal{M},$
$\quad\nabla\phi(\nabla\phi^{*}(y))=y,\quad\forall y\in\mathbb{R}^{d}.$

### v2

@@ Abstract
This article considers the generative modeling of the (mixed) states of quantum systems, and an approach based on denoising diffusion model is proposed. The key contribution is an algorithmic innovation that respects the physical nature of quantum states. More precisely, the commonly used density matrix representation of mixed-state has to be complex-valued Hermitian, positive semi-definite, and trace one. Generic diffusion models, or other generative methods, may not be able to generate data that strictly satisfy these structural constraints, even if all training data do. To develop a machine learning algorithm that has physics hard-wired in, we leverage mirror diffusion and borrow the physical notion of von Neumann entropy to design a new map, for enabling strict structure-preserving generation. Both unconditional generation and conditional generation via classifier-free guidance are experimentally demonstrated efficacious, the latter enabling the design of new quantum states when generated on unseen labels.
1 Introduction
Generative modeling, a specialized subfield of machine learning, is dedicated to creating models capable of generating new data points that closely match the statistical distribution of a given dataset. These models adeptly capture and replicate the patterns, structures, and characteristics of the input data, allowing them to yield new instances that seem to originate from the same distribution. Among them, diffusion model [50, 24, 51] has risen to prominence for its remarkable ability to produce high-quality data, especially notable in image generation [49, 25]. The fundamental mechanism of diffusion model starts with adding noise to the data and progressively altering it into a tractable Gaussian prior with analytically available transition kernels. The model then effectively reverses this noise addition, reconstructing the data from its noised state through the application of learned score functions [1, 22]. This process not only results in easily obtainable conditional score functions but also simplifies regression objectives, significantly boosting their capability in managing complex, high-dimensional data distributions.
Diffusion model has showcased impressive capabilities in generating various types of dataset across different application area, such as video synthesis [20, 4], language modeling [2, 41] and point cloud generation [60, 44]. With the recent rise of AI4Science, the potential of diffusion model in tackling scientific problems also started being explored, including applications in biology [31, 59], chemistry [15, 27], and climate science [45]. Other important notions relevant to this article include constrained generations [e.g., 39, 19, 18, 40] and conditional generations [e.g., 13, 26].
@@ 1 Introduction
What is under-studied but important is the (classical) generative modeling of quantum data. This is not the same as quantum computing or quantum machine learning, because the goal is not to design quantum algorithms, but to develop algorithms that run on classical computers for handling quantum data.
Similar to the already extremely successful application of the generation (and thus design) of molecular configurations [e.g., 58, 15], it would be useful to generate, for example, the states of quantum systems, before quantum computers become accessible to general researchers.
@@
To be able to represent quantum states, which in general can be mixed states, one needs to be mindful of their structural constraints that originated from physics. For example, it is standard in quantum information and quantum computing to represent mixed states as complex-valued density matrices [47], but such matrices have to be Hermitian, positive semidefinite, and trace one. Data-based generation of quantum density matrices therefore amounts to generative modeling in non-Euclidean spaces. Without strictly satisfying the structural constraints, the generated matrices will no longer correspond to quantum states [e.g., 28]. Unfortunately, generic generative modeling approaches will almost surely violate the constraints due to indispensable errors originated from finite data and compute. This article proposes the first machine learning methodology that generates density matrices similar to training data and strictly satisfying their structural constraints, based on hard-wiring these constraints into diffusion models via technique from convex optimization.
@@
Imperative to mention is, machine learning has already been demonstrated promising for quantum problems, and it is rapidly becoming popular in the domain. There is even an interesting research frontier on quantum algorithms for generative modeling [e.g., 11, 34, 7, 52, 61, 8]; however, this article still focuses on classical algorithm as quantum computer is not yet widely available, and existing results did not generate mixed states larger than 3-qubits. Meanwhile, classical algorithms have already been considered for many important quantum problems, such as the classification and regression of quantum states [e.g., 28, 29] and quantum state tomography [e.g., 54, 55, 6, 56, 30]. The latter, roughly speaking, corresponds to the process of estimating one latent quantum density matrix through repeated, but necessarily noisy measurements of many copies of the same state. It is very important to clarify the similarity and difference between this work and quantum state tomography:
@@
In fact, some may feel that quantum state tomography is already generative modeling, in a broad sense, because once the latent deterministic density matrix is recovered from noisy measurements, one can produce additional, synthetic classical measurements, which are random and ideally identically distributed with the training data. However, it is actually very different from the generative modeling of quantum states considered here, which might be in a narrow sense, however being a common setup in (classical) machine learning. More precisely, what this article considers is a collection of many different quantum mixed states, instead of classical observations of a single quantum state, and the goal is to generate more mixed states that are similar to the given ones. The problem of using machine learning to infer a single quantum state from classical observational data, i.e. state tomography, has already been considered by a series of seminal work mentioned above; there the probability distribution of data is uniquely determined by a latent quantum state, and thus one essentially needs to learn a $D$-dimensional vector (if the latent state is an $N$-qubit mixed state represented by a $2^{N}\times 2^{N}$ density matrix, then $D=2^{2N}$). The new task of state generative modeling in this article, on the other hand, assumes mixed states are already given and they follow some unknown latent probability distribution, which can be represented by a probability density function of $D$ variables, and the goal is in some sense to learn this function, which is exponentially more complex than a $D$-dimensional vector. One extra difference between this work and existing milestones on generative modeling for state tomography is, from a machine learning methodology point of view, this work is based on diffusion for generative modeling, while work cited above used, e.g., EBM, RBM, RNN instead.
@@
One perspective is to actually create innovative states. Unconditional generation will simply produce states similar to those in the training data, but modern (diffusion) generative modeling technique also unleashes the ability of conditional generation. More precisely, given training data that carry different labels (e.g., traditionally labels could be ‘astronaut’, ‘horse’, etc., but in our case they can be ‘insulating’, ‘magnetic’, ‘topological’, etc.), we can not only generate under a label that already has training data (e.g., ‘astronaut’, or ‘magnetic’), but also generate under a label that has no training data (e.g., the famous ‘horse-riding astronaut’ [23], or, this time, a motivating perspective of creating ‘topological magnetic’ states). This article humbly starts a first step toward this goal of scientific innovation, essentially via the interpolation/extrapolation of quantum data. The inter-/extra-polation, however, is in the sense of probability distributions but not sample points, because when one tries to design new states, it may not make sense to take, e.g., a specific ‘topological’ state and a specific ‘magnetic’ state and force a new ‘topological magnetic’ state based on them two alone. Instead, our conditional generative modeling approach hybrids multiple probability distributions and innovate just like prevailing generative AI tools.
@@
Another application is, generative modeling can also estimate the probability distribution of states in the data, hence providing further quantitative insights for probing the physical process behind it. Some consideration of this aspect was already included in the state tomography literature [e.g., 6], but not for an ensemble of states. Moreover, diffusion model excels in this regard. As a more recent generative approach, it is posited to be particularly effective for estimating density near the data manifold, even when data is high-dimensional but sample size is small, as long as there is a low-dimensional structure in the data. This is oftentimes the case for image and video generations, and their empirical successes were the root of this belief. Quantum density matrices data could be similar; at least, they are high-dimensional since dimension grows exponentially with the number of qubits, but scarce due to the cost of experiments. This is a case where diffusion generative model could be advantageous over traditional approaches.
@@
In addition, generative model can serve as a sampler for crafting synthetic data that are experimentally expensive to harvest. Synthetic data that are statistically similar to training data are useful for discovering system properties through helping estimate observable values. The seminal work [6] showcases one such application for state-tomography type generative modeling. The idea is, generative modeling could help obtain more virtual ‘measurements’ from physical ones, hence improving the accuracy of the estimated latent state. Analogously, generative modeling of a distribution of mixed states can help the investigation of stochastic quantum systems, such as Random Transverse-field Ising Models (RTFIM) [16, 17, 37] and measurement-altered quantum critical systems [46]. The generated states can empower more accurate estimation of random parameters in the system, such as coupling and transverse field strengths.
@@ 2.1 Unconditional Generation via Diffusion Model
Denoising diffusion generative models in Euclidean space admit many descriptions, with focuses on different perspectives, and here we adopt the stochastic differential equations (SDE) description [51]. Given samples of $\mathbb{R}^{d}$-valued random variable $x_{0}$ that follows the data distribution $p_{0}$, denoising diffusion adopts a forward noising process followed by a backward denoising generation process to generate more samples of $p_{0}$.
@@
The backward process then utilizes the time-reversal of the SDE (1) [1]. More precisely, if one considers
$\mathrm{d}y_{\tau}$ $=(-f(y_{\tau},T-\tau)+g(T-\tau)^{2}s(y_{\tau},T-\tau))\mathrm{d}\tau+g(T-\tau)\mathrm{d}w_{\tau},$ (2)
with $y_{0}\sim p_{T}$, where $s$, known as the score function, is given by $s(x,t)=\nabla\log p_{t}(x)$, then we have $y_{\tau}\sim p_{T-\tau}$, i.e. $y_{\tau}=x_{T-\tau}$ in distribution. In particular, the $T$-time evolution of (2), $y_{T}$, will follow the data distribution $p_{0}$.
@@
To do so, the score $s(x,t)=\nabla\log p_{t}(x)$ is often approximated using a neural network $s_{\theta}(x,t)$. For linear forward SDE, it is typically trained by minimizing an objective based on denoising score matching [57], namely
@@
where $\nabla\log p_{t}(x_{t}|x_{0})$ is the conditional score derived from (1) with a given initial condition.
@@ 2.2 Diffusion Guidance for Conditional Generation
Dhariwal and Nichol [13] exploited this fact and proposed the technique of classifier-based diffusion guidance, originally for boosting the sample quality generated by diffusion models. Instead of $p(x|y)$, they proposed to sample from a condition-enhanced distribution $p_{\gamma}(x|y)\propto p(x)p(y|x)^{\gamma}$. The score of this condition-enhanced distribution can be computed by
@@
$\gamma$ is called the strength of the guidance, which amplifies the influence of the conditioning when setting to a scale larger than $1$. When $\gamma>1$, the distribution is sharpened and focused onto its mode that corresponds to the condition $y$ [14]. By tuning $\gamma$, classifier guidance allows us to capture the influence of the condition signal.
However, classifier-based diffusion guidance can become impractical due to the expensive cost of training a separate classifier model. Such a classifier often requires to be noise-robust to handle the noise-corrupted input $x$ during the sampling process. [26] proposed classifier-free diffusion guidance to sample from the condition-enhanced distribution $p_{\gamma}(x|y)$ without the need for an extra discriminative model. Their insight came from another writing of Bayes’ rule: $\nabla\log p_{\gamma}(x|y)$ can also be decomposed as,
@@ 2.3 Density Matrix Representation of Quantum State
Precisely, a (quantum) density matrix is a matrix with each element being a complex number, which satisfies the following structural constraints [e.g., 47]:
@@
A $n-$qubit quantum state $\rho$ with full rank is an element of $\mathbb{C}^{+}_{2^{n}}$ with trace 1.
@@ 3 Method
Our task is to generate new samples of quantum mixed states, from a distribution characterized by the training data. Out of the three structure constraints of density matrix representation of quantum states, Hermitianity is easy to handle as one can just work with the upper triangular part of the matrix. Trace one condition will only be minimally violated due to generation error, and we simply normalize the generated result by dividing its trace to strictly enforce this constraint. The most nontrivial constraint is the positive semi-definiteness. Since this is a convex constraint, we can leverage the recent approach of Mirror Diffusion Model (MDM) [39], which is a class of diffusion models that generate data on convex-constrained sets, to enable exact generation on $\mathbb{C}^{+}_{n}$. To achieve this goal, we work out a complex and multivariate extension of the results in MDM.
@@ 3.1 Mirror Map with Real Numbers
Its inverse enjoys a pleasant property from convex analysis [48]: by defining the dual of $\phi$ as $\phi^{*}(y)=\sup_{x\in\mathcal{M}}\langle x,y\rangle-\phi(x),$ the inverse (as mapping) of $\nabla\phi$ is the same as the gradient of $\phi^{*}$, i.e.
$\nabla\phi^{*}(\nabla\phi(x))=x,\ \forall x\in\mathcal{M},\quad\nabla\phi(\nabla\phi^{*}(y))=y,\ \forall y\in\mathbb{R}^{d}.$